OpenStudy (trewolf):

Ron gets $5 on the first day at work, the square of that amount on the second day, and three times the amount of the second day on the third day. How much does Ron earn in three days?

OpenStudy (calculusxy):

First Day : $5 Second Day: Square of $5, which is \(5^2 = \) Third Day: Three times more than the square of $5, which is \(5^2 \times 3 = \) So evaluate: \(5 + 5^2 + (5^2 \times 3)\)

OpenStudy (trewolf):

A$45 B$75 C$105 D$125

OpenStudy (calculusxy):

Did you solve \(\large 5 + 5^2 + (5^2 \times 3)\)

OpenStudy (trewolf):

No im confused still @.@

OpenStudy (trewolf):

@calculusxy

OpenStudy (calculusxy):

Where are you exactly confused at?

OpenStudy (trewolf):

i don't know solving it

OpenStudy (trewolf):

@calculusxy

OpenStudy (calculusxy):

so on the first day, we know that he started off with $5.

OpenStudy (calculusxy):

on the second day, we know that rob squared the amount he got on the first day (which is $5). square is the same thing as: 5 x 5 or \(5^2\). you got that?

OpenStudy (trewolf):

5x5 = 25 @calculusxy

OpenStudy (trewolf):

5+5=10+25?

OpenStudy (calculusxy):

correct. on the third day, it's 3 times the square of 5, which means \(3 \times 5^2\).

OpenStudy (calculusxy):

you don't need to add the extra $5. he started it off with $5 so it would just be a single $5.

OpenStudy (trewolf):

75?

OpenStudy (trewolf):

@calculusxy

OpenStudy (calculusxy):

correct. now you need to add all of the amounts for all the three days. \(5 + 25 + 75\)

OpenStudy (trewolf):

105 @calculusxy

OpenStudy (calculusxy):

correct